OpenStudy (anonymous):

Can you tell me how to factor the quadratic for 7x^2+19x-6

OpenStudy (anonymous):

alright ill help

OpenStudy (anonymous):

Looking at this quadratic you have a=7 b=19 and c =-6

OpenStudy (anonymous):

so 7 times -6 is?...-42

OpenStudy (anonymous):

you need to find factors of -42 that add up to 19

OpenStudy (anonymous):

are you following?

OpenStudy (anonymous):

Yes! THis is very helpful

OpenStudy (anonymous):

your welcome. what do you think your factors will be?

OpenStudy (anonymous):

Since -42 is Negative one factor has to be positive and other one be negative REMEMBER this!

OpenStudy (anonymous):

2,3,7

OpenStudy (anonymous):

Your answer will be written similar to this form ..(x+1) (x+2)

OpenStudy (anonymous):

your coefficients are 7,19&-6 remember

OpenStudy (anonymous):

I know the form, i really just get stuck on the factoring

OpenStudy (anonymous):

What do you have so far

OpenStudy (anonymous):

im stuck on getting the factors of 42 that add up to 19

OpenStudy (anonymous):

Do you know what a factor tree or table is?

OpenStudy (anonymous):

i have an idea,

OpenStudy (anonymous):

Well factor down -42

OpenStudy (anonymous):

Have you checked your textbook on methods to solve this? You need to read up on factoring, because i can't just give you the answer. for this polynomial form ax^2+bx+c

OpenStudy (anonymous):

Ok done. I have the factors how do i out them in (x-1)(x+1) format?

OpenStudy (anonymous):

2/7 x -3 = -6/7 which is c/a and 2/7-3=-19/7 which is b/a so insert 2/7 as the right hand term of one factor and -3 as the right hand term of the other factor.. (x+2/7)(x-3).... are u getting this?

OpenStudy (anonymous):

Do you know how to remove the fraction?

OpenStudy (anonymous):

You can do it i promise!

OpenStudy (anonymous):

Multiply the first term of the factor by the denominator of the second term.

OpenStudy (anonymous):

This is the format of my math homework

OpenStudy (anonymous):

Your going to need to do some pencil & paper work hun.

OpenStudy (anonymous):

Multiply and you will be done!!!!!!

OpenStudy (anonymous):

I can't just give you the answer.. you wont learn anything.

OpenStudy (anonymous):

(x+2/7) (x-3)

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

ok... well you need to multiply by 7...

OpenStudy (anonymous):

cmon

OpenStudy (anonymous):

multiply 7 by 3?

OpenStudy (anonymous):

noo. leave (x-3) alone.

OpenStudy (anonymous):

(7x+2)(x-3) is your answer...

OpenStudy (anonymous):

You need to ask an instructor for help. Read up on your vocab and methods to help.

OpenStudy (anonymous):

Thanks for your help. Im knew to math.

OpenStudy (anonymous):

what grade are you in?